OpenStudy (anonymous):

if sin x=.21 what is cos x?

OpenStudy (anonymous):

use sinx^x+cosx^+1=0

OpenStudy (anonymous):

\[sinx^2+cosx^2+1=0\]

OpenStudy (anonymous):

this is hard..im bad at trig

OpenStudy (anonymous):

i dont get it..

OpenStudy (anonymous):

\[cosx^2=1-sinx^2\] \[cosx^2=1-(0.21)^2\] \[cosx^2=1-0.0441\] \[cosx^2=0.9559\] take square root both sides \[cosx=0.977\]

OpenStudy (anonymous):

do u understandd? if yes please click on best answer.

OpenStudy (anonymous):

thankks:)

OpenStudy (anonymous):

thankks:)

OpenStudy (anonymous):

u r welcome always.